# Graph Classification Gaussian Processes via Hodgelet Spectral Features

### v3

@@ 2 Gaussian Processes for Graph Classification
We assume a dataset containing $M$ undirected graphs $\mathcal{G}^{(1)},\dots,\mathcal{G}^{(M)}$ and labels $y^{(i)}\in\mathbb{Z}$, for $1\leq i\leq M$. We assign an orientation to each graph but emphasise that this choice is arbitrary. Let $\mathcal{V}^{(i)}$ be the vertex set and $\mathcal{E}^{(i)}$ be the edge set, both finite. For each graph $\mathcal{G}^{(i)}=(\mathcal{V}^{(i)},\mathcal{E}^{(i)})$, there are $N_{v}^{(i)}$ vertices, $N_{e}^{(i)}$ edges, and an incidence matrix $\boldsymbol{B}_{ve}^{(i)}\in\mathbb{Z}^{N_{v}^{(i)}\times N_{e}^{(i)}}$. The latter encodes the incidence between each vertex and edge, and furthermore defines the graph Laplacian $\boldsymbol{L}_{v}^{(i)}\coloneq\boldsymbol{B}_{ve}^{(i)}\boldsymbol{B}_{ve}^{(i)\top}\in\mathbb{Z}^{N_{v}^{(i)}\times N_{v}^{(i)}}$. By considering $3$-cliques, we obtain the edge-to-triangle incidence matrix $\boldsymbol{B}_{et}^{(i)}\in\mathbb{Z}^{N_{e}^{(i)}\times N_{t}^{(i)}}$, where $N_{t}^{(i)}$ is the number of triangles in $\mathcal{G}^{(i)}$. Likewise, we define the graph Helmholtzian $\boldsymbol{L}_{e}^{(i)}\coloneq\boldsymbol{B}_{ve}^{(i)\top}\boldsymbol{B}_{ve}^{(i)}+\boldsymbol{B}_{et}^{(i)}\boldsymbol{B}_{et}^{(i)\top}\in\mathbb{Z}^{N_{e}^{(i)}\times N_{e}^{(i)}}$, which applies to edges rather than vertices. We underline that $\boldsymbol{L}_{v}^{(i)}$ and $\boldsymbol{L}_{e}^{(i)}$ are instances of the discrete Hodge Laplacian (see Appendix D.2). Let $\mathcal{V}^{(i)}\rightarrow\mathbb{R}^{D_{v}}$ and $\mathcal{E}^{(i)}\rightarrow\mathbb{R}^{D_{e}}$ be two functions, for $D_{v},D_{e}\in\mathbb{N}$. By introducing an ordering on vertices and edges, we represent the preceding functions as matrices $\mathbb{R}^{D_{v}\times N_{v}^{(i)}}$ and $\mathbb{R}^{D_{e}\times N_{e}^{(i)}}$, respectively, which are understood as vertex and edge feature matrices, containing $D_{v}$ and $D_{e}$ channels, respectively. We denote vertex features for channel $1\leq d\leq D_{v}$ by the column vector $\boldsymbol{x}^{(i)}_{vd}\in\mathbb{R}^{N_{v}^{(i)}}$ and edge features for channel $1\leq d\leq D_{e}$ by the column vector $\boldsymbol{x}^{(i)}_{ed}\in\mathbb{R}^{N_{e}^{(i)}}$. We stress that our approach can adapt to graphs that may have vertex features, edge features, or both.
@@ 2.3 Hodgelet spectral features
We generate graph spectral features from graph wavelet coefficients and then use them in downstream GP classification tasks (see Appendix A). By combining the wavelet transform (5) and the Hodge decomposition (6), we compute the wavelet coefficients $\hat{\boldsymbol{x}}^{(i)}_{vdjc},\hat{\boldsymbol{x}}^{(i)}_{vdjh}$ and $\hat{\boldsymbol{x}}^{(i)}_{edje},\hat{\boldsymbol{x}}^{(i)}_{edjc},\hat{\boldsymbol{x}}^{(i)}_{edjh}$ (see Appendix C for more details). We derive our Hodgelet spectral features by concatenating the 2-norm of the preceding wavelet coefficients across each wavelet filter and channel, resulting in the column vectors $\boldsymbol{v}^{(i)}_{c},\boldsymbol{v}^{(i)}_{h}\in\mathbb{R}^{W_{v}D_{v}}$ and $\boldsymbol{e}^{(i)}_{e},\boldsymbol{e}^{(i)}_{c},\boldsymbol{e}^{(i)}_{h}\in\mathbb{R}^{W_{e}D_{e}}$. These spectral representations, which are invariant to graph isomorphism, are then fed to our additive Hodgelet kernel
@@ Appendix C Hodgelet spectral features
where $e,c,h$ in the last entry of the subscripts indicate exact, co-exact and harmonic components, respectively. We also recall that the indices $d,j$ correspond to the channel and filter dimensions, respectively. Concretely, the components are computed as
@@ D.2 Hodge Laplacians
If one takes a step back, it becomes apparent that graphs and simplicial complexes have a chain-like structure: vertices are connected to edges and edges are connected to triangles. More precisely, we say that vertices have edges as their upper neighbours, edges have vertices as their lower neighbours and triangles as their upper neighbours, and triangles have edges as their lower neighbours. This leads us to the definition of two incidence matrices, $\boldsymbol{B}_{ve}\in\mathbb{Z}^{N_{v}\times N_{e}}$ and $\boldsymbol{B}_{et}\in\mathbb{Z}^{N_{e}\times N_{t}}$, sometimes called vertex-to-edge and edge-to-triangle incidence matrices, respectively. While $\boldsymbol{B}_{ve}$ is the incidence matrix between vertices and edges, $\boldsymbol{B}_{et}$ is the incidence matrix between edges and triangles. In a more general setting, these matrices are simply called boundary matrices and they are denoted by $\boldsymbol{B}_{1}$ and $\boldsymbol{B}_{2}$, respectively, since vertices are $1$-simplices and edges are $2$-simplices. This results in the general definition of a boundary matrix $\boldsymbol{B}_{k}\in\mathbb{Z}^{N_{k}\times N_{k+1}}$, where $N_{k}$ is the number of $k$-simplices. Note that for a simplicial $K$-complex, $\boldsymbol{B}_{0}=\boldsymbol{0}$ and $\boldsymbol{B}_{k+1}=\boldsymbol{0}$. This simply means that 00-simplices have no lower neighbours, and $k$-simplices have no upper neighbours. We can define a more general notion of Laplacian, called the Hodge Laplacian,
@@
A $k$-clique is a subset of $k$ vertices, such that every $k$ distinct vertices in the clique are adjacent. Intuitively, $2$-cliques are like considering triangles. Edges have no upper neighbours in simplicial $1$-complexes, i.e. graphs. Hence, the edge-to-triangle incidence matrix $\boldsymbol{B}_{2}=\boldsymbol{0}$. In turn, this implies that the graph Helmholtzian is equal to its lower Laplacian, i.e. $\boldsymbol{L}_{1}=\boldsymbol{B}_{1}^{\top}\boldsymbol{B}_{1}$. Nevertheless, we can consider the triangles of a graph in order to have a non-zero $\boldsymbol{B}_{2}$ (see Figure D.2). We can then have a complete graph Helmholtzian, i.e. $\boldsymbol{L}_{1}=\boldsymbol{B}_{1}^{\top}\boldsymbol{B}_{1}+\boldsymbol{B}_{2}\boldsymbol{B}_{2}^{\top}$.
Figure D.2: A graph is augmented by taking its $2$-cliques and then assigned an orientation.
@@ D.3 Hodge decomposition
The Helmholtz decomposition is often called the fundamental theorem of vector calculus. This famous theorem states that any 2D vector fields on a Euclidean domain can be expressed as the sum of two orthogonal components: (1) one that intuitively has no sinks or sources, called divergence-free or solenoidal, and (2) one that has no vortices, sometimes called curl-free or irrotational. This decomposition extends to differential forms on manifolds and is referred to as the Hodge decomposition. For Riemannian manifolds, we can identify differential $1$-forms with vector fields via the musical isomorphism. This implies that Hodge decomposition can be interpreted as a decomposition of vector fields on Riemannian manifolds. The Hodge decomposition on the $k$-simplex feature space $\mathbb{R}^{N_{k}}$ (for a single channel) leads to the sum of three orthogonal components: exact, co-exact, and harmonic,
@@ E.2 Vector Field Classification Details
We provide some details about the vector field classification experiment in Section 3. To generate a random vector field, we first sample a Gaussian process $f:\Omega\times\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$, and then take its derivatives $f_{x},f_{y}$ in both spatial components. Noting that a curl-free 2D vector field is always the gradient of a potential, we can sample an arbitrary curl-free field by taking
@@
For the vector field $f$, we used samples of the squared-exponential Gaussian process, sampled using its random feature approximation [32]. We display this in Figure E.1, where we plot the derivatives $f_{x},f_{y}$, and their combinations to yield divergence-free and curl-free fields.
@@
To project a vector field $\boldsymbol{X}:\mathbb{R}^{2}\rightarrow\mathbb{R}$ onto the edges of a graph, we employ the de Rham map [33], which “discretises” a vector field into edge signals. For a given oriented edge $e=(v_{0},v_{1})\in\mathcal{E}$ with endpoint coordinates $\boldsymbol{x}_{0}$ and $\boldsymbol{x}_{1}$, we define the projection $X^{e}$ of $\boldsymbol{X}$ onto $e$ by the following integral,

### v4

@@ 2 Gaussian Processes for Graph Classification
We assume a dataset containing $M$ undirected graphs $\mathcal{G}^{(1)},\dots,\mathcal{G}^{(M)}$ and labels $y^{(i)}\in\mathbb{Z}$, for $1\leq i\leq M$. We assign an orientation to each graph but emphasise that this choice is arbitrary. Let $\mathcal{V}^{(i)}$ be the vertex set and $\mathcal{E}^{(i)}$ be the edge set, both finite. For each graph $\mathcal{G}^{(i)}=(\mathcal{V}^{(i)},\mathcal{E}^{(i)})$, there are $N_{v}^{(i)}$ vertices, $N_{e}^{(i)}$ edges, and an incidence matrix $\boldsymbol{B}_{ve}^{(i)}\in\mathbb{Z}^{N_{v}^{(i)}\times N_{e}^{(i)}}$. The latter encodes the incidence between each vertex and edge, and furthermore defines the graph Laplacian $\boldsymbol{L}_{v}^{(i)}\coloneq\boldsymbol{B}_{ve}^{(i)}\boldsymbol{B}_{ve}^{(i)\top}\in\mathbb{Z}^{N_{v}^{(i)}\times N_{v}^{(i)}}$. By considering $3$-cliques, we obtain the edge-to-triangle incidence matrix $\boldsymbol{B}_{et}^{(i)}\in\mathbb{Z}^{N_{e}^{(i)}\times N_{t}^{(i)}}$, where $N_{t}^{(i)}$ is the number of triangles in $\mathcal{G}^{(i)}$. Likewise, we define the graph Helmholtzian $\boldsymbol{L}_{e}^{(i)}\coloneq\boldsymbol{B}_{ve}^{(i)\top}\boldsymbol{B}_{ve}^{(i)}+\boldsymbol{B}_{et}^{(i)}\boldsymbol{B}_{et}^{(i)\top}\in\mathbb{Z}^{N_{e}^{(i)}\times N_{e}^{(i)}}$, which applies to edges rather than vertices. We underline that $\boldsymbol{L}_{v}^{(i)}$ and $\boldsymbol{L}_{e}^{(i)}$ are instances of the discrete Hodge Laplacian (see Appendix D.2). Let $\mathcal{V}^{(i)}\rightarrow\mathbb{R}^{D_{v}}$ and $\mathcal{E}^{(i)}\rightarrow\mathbb{R}^{D_{e}}$ be two functions, for $D_{v},D_{e}\in\mathbb{N}$. By introducing an ordering on vertices and edges, we represent the preceding functions as matrices $\mathbb{R}^{D_{v}\times N_{v}^{(i)}}$ and $\mathbb{R}^{D_{e}\times N_{e}^{(i)}}$, respectively, which are understood as vertex and edge feature matrices, containing $D_{v}$ and $D_{e}$ dimensions, respectively. We denote vertex features for dimension $1\leq d\leq D_{v}$ by the column vector $\boldsymbol{x}^{(i)}_{vd}\in\mathbb{R}^{N_{v}^{(i)}}$ and edge features for dimension $1\leq d\leq D_{e}$ by the column vector $\boldsymbol{x}^{(i)}_{ed}\in\mathbb{R}^{N_{e}^{(i)}}$. We stress that our approach can adapt to graphs that may have vertex features, edge features, or both.
@@ 2.3 Hodgelet spectral features
We generate graph spectral features from graph wavelet coefficients and then use them in downstream GP classification tasks (see Appendix A). By combining the wavelet transform (5) and the Hodge decomposition (6), we compute the wavelet coefficients $\hat{\boldsymbol{x}}^{(i)}_{vdjc},\hat{\boldsymbol{x}}^{(i)}_{vdjh}$ and $\hat{\boldsymbol{x}}^{(i)}_{edje},\hat{\boldsymbol{x}}^{(i)}_{edjc},\hat{\boldsymbol{x}}^{(i)}_{edjh}$ (see Appendix C for more details). We derive our Hodgelet spectral features by concatenating the 2-norm of the preceding wavelet coefficients across each wavelet filter and dimension, resulting in the column vectors $\boldsymbol{v}^{(i)}_{c},\boldsymbol{v}^{(i)}_{h}\in\mathbb{R}^{W_{v}D_{v}}$ and $\boldsymbol{e}^{(i)}_{e},\boldsymbol{e}^{(i)}_{c},\boldsymbol{e}^{(i)}_{h}\in\mathbb{R}^{W_{e}D_{e}}$. These spectral representations, which are invariant to graph isomorphism, are then fed to our additive Hodgelet kernel
@@ Appendix C Hodgelet spectral features
where $e,c,h$ in the last entry of the subscripts indicate exact, co-exact and harmonic components, respectively. We also recall that the indices $d,j$ correspond to the dimension and filter dimensions, respectively. Concretely, the components are computed as
@@ D.2 Hodge Laplacians
If one takes a step back, it becomes apparent that graphs and simplicial complexes have a chain-like structure: vertices are connected to edges and edges are connected to triangles. More precisely, we say that vertices have edges as their upper neighbours, edges have vertices as their lower neighbours and triangles as their upper neighbours, and triangles have edges as their lower neighbours. This leads us to the definition of two incidence matrices, $\boldsymbol{B}_{ve}\in\mathbb{Z}^{N_{v}\times N_{e}}$ and $\boldsymbol{B}_{et}\in\mathbb{Z}^{N_{e}\times N_{t}}$, sometimes called vertex-to-edge and edge-to-triangle incidence matrices, respectively. While $\boldsymbol{B}_{ve}$ is the incidence matrix between vertices and edges, $\boldsymbol{B}_{et}$ is the incidence matrix between edges and triangles. In a more general setting, these matrices are simply called boundary matrices and they are denoted by $\boldsymbol{B}_{1}$ and $\boldsymbol{B}_{2}$, respectively, since vertices are 00-simplices and edges are $1$-simplices. This results in the general definition of a boundary matrix $\boldsymbol{B}_{k}\in\mathbb{Z}^{N_{k}\times N_{k+1}}$, where $N_{k}$ is the number of $k$-simplices. Note that for a simplicial $K$-complex, $\boldsymbol{B}_{0}=\boldsymbol{0}$ and $\boldsymbol{B}_{k+1}=\boldsymbol{0}$. This simply means that 00-simplices have no lower neighbours, and $k$-simplices have no upper neighbours. We can define a more general notion of Laplacian, called the Hodge Laplacian,
@@
A $k$-clique is a subset of $k$ vertices, such that every $k$ distinct vertices in the clique are adjacent. Intuitively, $3$-cliques are like considering triangles. Edges have no upper neighbours in simplicial $1$-complexes, i.e. graphs. Hence, the edge-to-triangle incidence matrix $\boldsymbol{B}_{2}=\boldsymbol{0}$. In turn, this implies that the graph Helmholtzian is equal to its lower Laplacian, i.e. $\boldsymbol{L}_{1}=\boldsymbol{B}_{1}^{\top}\boldsymbol{B}_{1}$. Nevertheless, we can consider the triangles of a graph in order to have a non-zero $\boldsymbol{B}_{2}$ (see Figure D.2). We can then have a complete graph Helmholtzian, i.e. $\boldsymbol{L}_{1}=\boldsymbol{B}_{1}^{\top}\boldsymbol{B}_{1}+\boldsymbol{B}_{2}\boldsymbol{B}_{2}^{\top}$.
Figure D.2: A graph is augmented by taking its $3$-cliques and then assigned an orientation.
@@ D.3 Hodge decomposition
The Helmholtz decomposition is often called the fundamental theorem of vector calculus. This famous theorem states that any 2D vector fields on a Euclidean domain can be expressed as the sum of two orthogonal components: (1) one that intuitively has no sinks or sources, called divergence-free or solenoidal, and (2) one that has no vortices, sometimes called curl-free or irrotational. This decomposition extends to differential forms on manifolds and is referred to as the Hodge decomposition. For Riemannian manifolds, we can identify differential $1$-forms with vector fields via the musical isomorphism. This implies that Hodge decomposition can be interpreted as a decomposition of vector fields on Riemannian manifolds. The Hodge decomposition on the $k$-simplex feature space $\mathbb{R}^{N_{k}}$ (for a single dimension) leads to the sum of three orthogonal components: exact, co-exact, and harmonic,
@@ E.2 Vector Field Classification Details
We provide some details about the vector field classification experiment in Section 3. To generate a random vector field, we first sample a Gaussian process $f:\Omega\times\mathbb{R}^{2}\rightarrow\mathbb{R}$, and then take its derivatives $f_{x},f_{y}$ in both spatial components. Noting that a curl-free 2D vector field is always the gradient of a potential, we can sample an arbitrary curl-free field by taking
@@
For the scalar field $f$, we used samples of the squared-exponential Gaussian process, sampled using its random feature approximation [32]. We display this in Figure E.1, where we plot the derivatives $f_{x},f_{y}$, and their combinations to yield divergence-free and curl-free fields.
@@
To project a vector field $\boldsymbol{X}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ onto the edges of a graph, we employ the de Rham map [33], which “discretises” a vector field into edge signals. For a given oriented edge $e=(v_{0},v_{1})\in\mathcal{E}$ with endpoint coordinates $\boldsymbol{x}_{0}$ and $\boldsymbol{x}_{1}$, we define the projection $X^{e}$ of $\boldsymbol{X}$ onto $e$ by the following integral,